OpenStudy (anonymous):

Can you help me please: Find the value of 3x/sqrt(x^3)?

OpenStudy (rajee_sam):

\[\frac{ 3x }{ \sqrt{x ^{3}} } = \frac{ 3x }{ x \sqrt{x} } = \frac{ 3 }{ \sqrt{x} }\]

OpenStudy (anonymous):

Thank you so much. I see where I went wrong.